OpenStudy (anonymous):

Find the solution of D.E Xe(x(2)+y) dx= y dy

OpenStudy (nenadmatematika):

you haven't written this properly...post again more carefuly :D

OpenStudy (anonymous):

Yeah, I can't read this.

OpenStudy (anonymous):

X Exp(x2+y) = y dy

OpenStudy (anonymous):

\[xe^{x^2+y} dx=ydy?\]

OpenStudy (anonymous):

sorry i'm new >> yeah that is right please help me

OpenStudy (anonymous):

That's okay, you will learn. Back to the DE in our hands. This DE is separable since we can write it as \(x(e^{x^2}e^y)dx=y dy.\). Divide both sides by \(e^y\) and then integrate both sides.

OpenStudy (anonymous):

Dividing both sides by \(e^y\) gives \[xe^{x^2}dx=\frac{y}{e^y}dy \implies \int xe^{x^2}dx=\int ye^{-y}dy.\]

OpenStudy (anonymous):

Do you know how to evaluate the integrals?

OpenStudy (anonymous):

what (frac)

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Hmm looks like your browser has some problems with showing latex. Anyways frac{x}{y} means x over y.

OpenStudy (anonymous):

what about Homogeneous ?

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

can solve with it ? is that Homogeneous Eq.?

OpenStudy (anonymous):

It doesn't look homogeneous to me.

OpenStudy (anonymous):

if it ? how i solve in this way? and can you help me to get me book with maths Eq.

OpenStudy (anonymous):

Are you asking about how to solve Homogeneous DE's?

OpenStudy (anonymous):

These are good notes on HDE http://tutorial.math.lamar.edu/Classes/DE/HOHomogeneousDE.aspx

OpenStudy (anonymous):

thanks a lot ..... <3

OpenStudy (anonymous):

Glad to help.